Report on τ → pℓℓ

Marcin Chrząszcz, Alberto Lusiani

Institute of Nuclear Physics, INFN, Scuola Normale

Superiore

27^{th} March 2013 MC & data

How does truth matching work?

Let’s say we what to study a decay of D^{0} → πK . 1 From π and D^{0} you construct the D^{0} with background. 2 At some point you may want to know how many actual D^{0} survived

the selection. Then you do in the code (or some module does it for you) mcFromReco(...).

3 However the number this function returns will be an

underestimation of the actual number. 4 If FSR happen you will end up with a channel D^{0} → πK γ, which

software will not match with search channel: D^{0} → πK .

How does truth matching work?

How ever code does one thing for you:

1 For example if you have a decay: τ^{−} → π^{+}π^{−}π^{−}ν_{τ} 2 Of coz there is no ν BtaCandidate list . 3 Same thing was implemented for FSR correction. 4 The new code should not create problems with the old modules

etc.